# Prithvi-EO-2.0: A Versatile Multi-Temporal Foundation Model for Earth Observation Applications

### v1

@@ 2 Dataset Description and Sampling
The HLS data [6] is the harmonized surface reflectance data product from NASA/USGS’s Landsat 8 and 9 and the ESA’s Sentinel-2A and Sentinel-2B satellites. The HLS data is compatible with the 40-year Landsat data record, but it has a temporal resolution of two-three days on average. Fifteen visible and infrared bands are available in HLS, however, not all channels are present in Sentinel and Landsat. Therefore, we used 6 channels common to both Sentinel and Landsat, namely 2, 3, 4, 8A, 11, and 12 for training our model.
@@
Our sampling approach aimed at creating a high-quality dataset representing diverse land use and ecosystems for robust model training and validation, while minimizing cloud and missing data issues. We obtained this with the following steps. First, we began by calculating the proportion of LULC classes and ecoregions for each HLS tile using the Copernicus Land Cover 100m [7] and RESOLVE Ecoregions [8] labels. Second, after merging the twelve closed and open forest classes into two classes, we sampled 100 tiles per LULC class from the 500 tiles with the highest class proportion. Urban areas are upsampled by selecting 1,000 tiles that cover about 60% of the global urban areas. Additionally, we included 1,000 tiles with a high LULC class entropy to capture heterogeneous landscapes. Next, we ensured that the 846 ecoregions were represented if the size of the region allowed for it. In total, 712 ecoregions are present in three or more tiles, while 68 are not specifically included due to limited area coverage. Using a 95%-5% train-validation split, this process resulted in 3,156 train and 168 validation tiles. 133 of these tiles are missing in the final dataset due to quality issues or insufficient data coverage, e.g., in Greenland and Antarctica. The LULC distribution of the final dataset is visualized in Figure 1, clearly showing that sea and bare vegetation (desert) are down-sampled compared to all 18k HLS land tiles. Because we up-sample build-up areas, managed vegetation is also over-represented due to the geographical proximity.
@@
Figure 2: Global HLS dataset distribution visualized on a tile-level. The number of training samples are color-coded in orange to green while validation tiles are visualized in blue.
@@ 3 Pretraining and Model architecture
Based on the above pretraining routine and architecture, we developed Prithvi-EO-2.0 in two sizes, 300M and 600M, based on ViT-L and ViT-H, respectively. We trained both versions with (Prithvi-EO-2.0$\ast$-TL) and without (Prithvi-EO-2.0$\ast$) temporal and location information. The models were trained for 400 epochs; the 300M models were trained on 80 GPUs (utilizing $\sim$21.000 GPU-hours for the full training), while the 600M models were on 240 GPUs (utilizing $\sim$58.000 GPU-hours for the full training). The local batch size (i.e., the batch size per GPU) for ViT-L was 48, and for ViT-H, was 16 to accommodate the larger model; a global batch size (i.e., the local batch size multiplied by the number of utilized GPUs) of 3,840 is then effectively used for all models. The weight decay was set to 0.05, and we used a cosine scheduler with a maximum learning rate of 5e-4, linear warmup for 40 epochs starting from a learning rate of 1e-6. For models with temporal and location data, we set the drop probability to 0.1. As described in section 2, the training samples have shape $4\times 256\times 256$ for time, height, and width, respectively, and we used random crop (final input shape is $4\times 224\times 224$) and random horizontal flips as data augmentation. Data parallelism in HPC systems is essential for efficiently handling large-scale remote sensing data required for training deep learning models [14]. By distributing the data across multiple GPUs, HPC systems can significantly reduce training time while maintaining accuracy. PyTorch DDP was used to scale the model pretraining. The models were pretrained on the Booster partition of the JUWELS (Jülich Wizard for European Leadership Science) supercomputer [^3] https://apps.fz-juelich.de/jsc/hps/juwels/index.html [15] hosted at JSC, which features 936 compute nodes, each equipped with 4 NVIDIA A100 GPUs.
@@ 4.1 Benchmarking
We used the most popular and rigorous benchmark framework available for Earth Observation foundation models: GEO-Bench [5]. GEO-Bench presents two key characteristics. First, it includes a carefully collected list of 6 classification and 6 semantic segmentation datasets for different resolutions and domains (see Table 1 and Figure 4). Second, it established a protocol that allows to compare different models fairly and reproducibly. This starts with hyperparameter tuning, where each model is allowed a budget of tries, 10 in our study, on each dataset. Then, the best hyperparameters found for each dataset on the validation set are used for repeated experiments (N=10) on each dataset with different seeds. This allows to account for randomness that is present when training AI models, and more confidently compare results.
@@
Figure 4: Examples of images and labels for the GEO-Bench datasets [5].
@@
Results of benchmarking are shown in Figures 6, 5 and 7. Figure 6 shows the average performance across the classification and segmentation datasets, while Figure 5 depicts the overall performance across all 12 datasets. These figures highlight that, for the classification tasks, DOFA [20] comes first, with Prithvi-EO-2.0-600M-TL close second and Prithvi-EO-2.0-600M third. For segmentation Prithvi-EO-2.0-600M-TL and Prithvi-EO-1.0-600M top the leaderboard with some margin on DeCUR [19] and DINO [18] ResNets. Overall, Prithvi-EO-2.0-600M-TL and Prithvi-EO-2.0-600M obtained the best combined performance across all 12 datasets. Note that in this case we are combining different metrics for different tasks (i.e., overall accuracy for classification and mean Jaccard Index for segmentation). However, such metrics are all between 0 and 1 and models are treated equally. Therefore, we believe this procedure is appropriate to get an overall performance picture across GEO-Bench datasets.
Figure 5: Overall average performance across all GEO-Bench datasets.
Figure 6: Averages for datasets over GEO-Bench classification (Overall accuracy) and segmentation (mean Jaccard Index) datasets.
Figures 5 and 6 are important to get an overall picture, while Figure 7 gives a more detailed view on the distribution of model’s performance across different datasets. Overall, it is interesting to see that some datasets carry very little insights to differentiate models, as performance is generally extremely high and condensed (e.g., m-brik-klin or m-pv4ger). For the other datasets, Prithvi-EO-2.0-600M-TL and Prithvi-EO-2.0-600M a the best performer in four out of six tasks targeting medium resolution (i.e. S2 at 10m), with tasks like m-bigearthnet and m-SA-crop-type classification that showed better performance by all Prithvi-EO-2.0-600M and Prithvi-EO-2.0-300M versions when compared to the other models. This is particularly interesting because these tasks focus on earth observation applications like land use classification using the CORINE Land Cover [22], which has strong focus on environmental monitoring categories (e.g. agriculture- and forest-related especially), and crop classification, which presents strong seasonality. These are exactly the type of domains that we targeted when crafting our multi-temporal global dataset, and these results suggest the model has learned relevant features to be highly performant on them.
@@
Figure 7: Distribution of performance across all models for classification (a) and segmentation (b) GEO-Bench [5] tasks in terms of overall accuracy and mean IoU, respectively. At the bottom of each graph, dataset input sensor and resolution.
@@ 4.2 Downstream tasks
We clustered the downstream tasks in three main areas: disaster response, land use and crop mapping, and multi-modal ecosystem dynamics monitoring. A full list can be found in Table 3
@@ 4.2.3 Multi-modal Ecosystem Dynamics
Above Ground Biomass (AGB) is a crucial climate variable that encompasses all living biomass above the soil, essential for understanding land use changes and carbon inventories. Although direct measurement of AGB is challenging, remote sensing methods and machine learning techniques can be used to estimate it, providing comprehensive spatial and temporal data. The BioMassters dataset aimed to explore deep learning approaches for predicting yearly Above Ground Biomass (AGB) in Finnish forests using multi-modal, multi-temporal Sentinel-1 and Sentinel-2 satellite data from 2016 to 20212. The dataset comprises 11,462 AGB reference images, each paired with 12 months of corresponding satellite imagery, covering 2,560$\times$2,560-meter forest patches, with AGB measured in metric tonnes per pixel and derived from LiDAR data calibrated with in-situ measurements. The winning model from the data challenge is used as the benchmark for comparison with the Prithvi model. This model is based on a U-Net architecture with temporal attention encoding [35], using all 12 time steps and all 15 channels from Sentinel-1 and Sentinel-2 imagery as inputs for each yearly AGB prediction. The winning model achieved an RMSE of 27.49 on the holdout test set.
@@
Photosynthesis (Gross Primary Productivity (GPP)), the critical process converting solar radiation into energy and materials that sustain life, represents the largest global carbon cycle flux, yet remains challenging to quantify accurately without direct observation [36]. Data-driven approaches using machine learning to integrate satellite data with site-level information (e.g. from eddy covariance measurement networks [37]) provide diagnostic constraints on global GPP, complementing process-based models [38]. However, these approaches face challenges due to scale mismatches between in-situ measurements and satellite data, as well as generalization issues, which could potentially be mitigated by using foundation models like Prithvi that embed generalized representations of high-resolution data [39].
This downstream application aims to evaluate the potential of Prithvi to estimate GPP at globally distributed eddy covariance sites. We utilize daily GPP data from eddy covariance flux towers, obtained from FLUXNET [40], AmeriFlux [^10] https://ameriflux.lbl.gov/data/flux-data-products/, and ICOS [41] networks. We use datasets processed by the ONEFLUX pipeline for high-quality and consistent data. The daily eddy covariance data includes individual files for each site. The target variable for analysis is “GPP_NT_VUT_REF", which corresponds to GPP estimates derived from Net Ecosystem Exchange (NEE) measurements using the nighttime approach. We selected daily GPP data with at least 60% of high-quality hourly or half-hourly data for temporal aggregation. Metadata for flux tower sites, including latitude and longitude, is recorded to note site locations where satellite-based regressor variables are collected to prepare the dataset.
@@
These steps are replicated with all four variants of Prithvi models and we quantify the model’s skill in predicting GPP using the coefficient of determination, $R^{2}$ measure between the predicted and true GPP values. The $R^{2}$ performance in each train test split across all years is shown in Figure 17. Prithvi is also compared to a baseline model based on random forest, which uses the spatial average of HLS bands across each chip and the MERRA-2 data, representing the state-of-the-art approaches [39]. The baseline model also incorporated vegetation indices, including EVI, Normalized Difference Vegetation Index (NDVI), Normalized Difference Water Index (NDWI), Green Chlorophyll Index (GCI), near-infrared reflectance of vegetation (NIRv), and kernel NDVI. Results in Table 13 show that Prithvi-EO-2.0 models performed better than the random forest without vegetation indexes with up-to 20% improvement on $R^{2}$. Furthermore, the Prithvi-EO-2.0-600M version with temporal and location embeddings consistently outperformed the random forest including the vegetation indexes despite not having this information in its input. The proposed fine-tuning approach using Prithvi-EO-2.0 and MERRA-2 data shows promise for expanding the study to additional flux tower sites and extrapolating GPP beyond the flux network to derive a key climatic variable.

### v2

@@ 2 Dataset Description and Sampling
The HLS data [6] is the harmonized surface reflectance data product from NASA/USGS’s Landsat 8 and 9 and the ESA’s Sentinel-2A and Sentinel-2B satellites. The HLS data is compatible with the 40-year Landsat data record, but it has a temporal resolution of two-three days on average. Fifteen visible and infrared bands are available in HLS, however, not all channels are present in Sentinel and Landsat. Therefore, we used six channels common to both, Sentinel and Landsat, namely Blue, Green, Red, NIR, SWIR1, and SWIR2.
@@
Our sampling approach aimed at creating a high-quality dataset representing diverse land use and ecosystems for robust model training, while minimizing cloud and missing data issues. We obtained this with the following steps. First, we began by calculating the proportion of LULC classes and ecoregions for each HLS tile using the Copernicus Land Cover 100m [7] and RESOLVE Ecoregions [8] labels. Second, after merging the 12 closed and open forest classes into two classes, we sampled 100 tiles per LULC class from the 500 tiles with the highest class proportion. Urban areas are upsampled by selecting 1,000 tiles that cover about 60% of the global urban areas. Additionally, we included 1,000 tiles with a high LULC class entropy to capture heterogeneous landscapes. Next, we ensured that the 846 ecoregions were represented if the size of the region allowed for it. In total, 712 ecoregions are present in three or more tiles, while 68 are not specifically included due to limited area coverage. We applied a 95%-5% train-validation split and dropped additional 133 sampled tiles due to quality issues or insufficient data coverage, e.g., in Greenland and Antarctica. This process resulted in 3,028 train and 163 validation tiles. The LULC distribution of the final dataset is visualized in Figure 1, clearly showing that sea and bare vegetation (desert) are down-sampled compared to all 18k HLS land tiles. Because we up-sampled build-up areas, managed vegetation is also over-represented due to the geographical proximity.
@@
Figure 2: Global HLS dataset distribution visualized on a tile-level. The number of training samples are color-coded in blue to green while validation tiles are visualized in magenta.
@@ 3 Pretraining and Model architecture
Based on the above pretraining routine and architecture, we developed Prithvi-EO-2.0 in two sizes, 300M and 600M, based on ViT-L and ViT-H, respectively. We trained both versions with (Prithvi-EO-2.0$\ast$-TL) and without (Prithvi-EO-2.0$\ast$) temporal and location information. The models were trained for 400 epochs; the 300M models were trained on 80 GPUs (utilizing $\sim$21.000 GPU-hours for the full training), while the 600M models were on 240 GPUs (utilizing $\sim$58.000 GPU-hours for the full training). The local batch size (i.e., the batch size per GPU) for ViT-L was 48, and for ViT-H, was 16 to accommodate the larger model; a global batch size (i.e., the local batch size multiplied by the number of utilized GPUs) of 3,840 is then effectively used for all models. The weight decay was set to 0.05, and we used a cosine scheduler with a maximum learning rate of 5e-4, linear warmup for 40 epochs starting from a learning rate of 1e-6. For models with temporal and location data, we set the drop probability to 0.1. As described in section 2, the training samples have shape $4\times 256\times 256$ for time, height, and width, respectively, and we used random crop (final input shape is $4\times 224\times 224$) and random horizontal flips as data augmentation. Data parallelism in HPC systems is essential for efficiently handling large-scale remote sensing data required for training deep learning models [14]. By distributing the data across multiple GPUs, HPC systems can significantly reduce training time while maintaining accuracy. PyTorch DDP was used to scale the model pretraining. The models were pretrained on the Booster partition of the JUWELS (Jülich Wizard for European Leadership Science) supercomputer [^3] https://apps.fz-juelich.de/jsc/hps/juwels/index.html [15] hosted at JSC, which features 936 compute nodes, each equipped with four NVIDIA A100 GPUs.
@@ 4.1 Benchmarking
We used the most popular and rigorous benchmark framework available for Earth Observation foundation models: GEO-Bench [5]. GEO-Bench presents two key characteristics. First, it includes a carefully collected list of six classification and six semantic segmentation datasets for different resolutions and domains (see Table 1 and Figure 4). Second, it established a protocol that allows to compare different models fairly and reproducibly. This starts with hyperparameter tuning, where each model is allowed a budget of tries, 10 in our study, on each dataset. Then, the best hyperparameters found for each dataset on the validation set are used for repeated experiments (N=10) on each dataset with different seeds. This allows to account for randomness that is present when training AI models, and more confidently compare results.
@@
Figure 4: (a) Classification and (b) segmentation examples in the GEO-Bench datasets [5].
@@
Results of benchmarking are shown in Figure 5, including the average performance across the classification and segmentation datasets, as well as the overall performance across all 12 datasets. These figures highlight that, for the classification tasks, DOFA [20] comes first, with Prithvi-EO-2.0-600M-TL close second and Prithvi-EO-2.0-600M third. For segmentation Prithvi-EO-2.0-600M-TL and Prithvi-EO-1.0-600M top the leaderboard with some margin on DeCUR [19] and DINO [18] ResNets. Overall, Prithvi-EO-2.0-600M-TL and Prithvi-EO-2.0-600M obtained the best combined performance across all 12 datasets. Note that in this case we are combining different metrics for different tasks (i.e., overall accuracy for classification and mean Jaccard Index for segmentation). However, such metrics are all between 0 and 1 and models are treated equally. Therefore, we believe this procedure is appropriate to get an overall performance picture across GEO-Bench datasets.
Figure 5: Aggregated performance across (a) all GEO-Bench datasets, (b) all classification tasks, and (c) all segmentation tasks. The Prithvi-EO-2.0 models are highlighted in blue.
Figure 5 presents the aggregated results, while Figure 6 gives a more detailed view on the distribution of model’s performance across different datasets. Overall, it is interesting to see that some datasets carry very little insights to differentiate models, as performance is generally extremely high and condensed (e.g., m-brik-klin or m-pv4ger). For the other datasets, Prithvi-EO-2.0-600M-TL and Prithvi-EO-2.0-600M a the best performer in four out of six tasks targeting medium resolution (i.e. S2 at 10m), with tasks like m-bigearthnet and m-SA-crop-type classification that showed better performance by all Prithvi-EO-2.0-600M and Prithvi-EO-2.0-300M versions when compared to the other models. This is particularly interesting because these tasks focus on earth observation applications like land use classification using the CORINE Land Cover [22], which has strong focus on environmental monitoring categories (e.g. agriculture- and forest-related especially), and crop classification, which presents strong seasonality. These are exactly the type of domains that we targeted when crafting our multi-temporal global dataset, and these results suggest the model has learned relevant features to be highly performant on them.
@@
Figure 6: Distribution of performance across all models from 10 repeated runs for the GEO-Bench (a) classification and (b) segmentation tasks in terms of accuracy and mean IoU, respectively. We list input sensor and resolution for each dataset.
@@ 4.2 Downstream tasks
We clustered the downstream tasks in three main areas: disaster response, land use and crop mapping, and multi-modal ecosystem dynamics monitoring. A full list can be found in Table 3.
@@ 4.2.3 Multi-modal Ecosystem Dynamics
Above Ground Biomass (AGB) is a crucial climate variable that encompasses all living biomass above the soil, essential for understanding land use changes and carbon inventories. Although direct measurement of AGB is challenging, remote sensing methods and machine learning techniques can be used to estimate it, providing comprehensive spatial and temporal data. The BioMassters dataset aimed to explore deep learning approaches for predicting yearly Above Ground Biomass (AGB) in Finnish forests using multi-modal, multi-temporal Sentinel-1 and Sentinel-2 satellite data from 2016 to 2022. The dataset comprises 11,462 AGB reference images, each paired with 12 months of corresponding satellite imagery, covering 2,560$\times$2,560-meter forest patches, with AGB measured in metric tonnes per pixel and derived from LiDAR data calibrated with in-situ measurements. The winning model from the data challenge is used as the benchmark for comparison with the Prithvi model. This model is based on a U-Net architecture with temporal attention encoding [35], using all 12 time steps and all 15 channels from Sentinel-1 and Sentinel-2 imagery as inputs for each yearly AGB prediction. The winning model achieved an RMSE of 27.49 on the holdout test set.
@@
Photosynthesis (Gross Primary Productivity (GPP)), the critical process converting solar radiation into energy and materials that sustain life, represents the largest global carbon cycle flux, yet remains challenging to quantify accurately without direct observation at regional to global scales [36]. Data-driven approaches using machine learning to integrate satellite data with site-level information (e.g. from eddy covariance measurement networks [37]) provide diagnostic constraints on global GPP, complementing process-based models [38]. However, these approaches face challenges due to scale mismatches between in-situ measurements and satellite data, as well as generalization issues, which could potentially be mitigated by using foundation models like Prithvi that embed generalized representations of high-resolution data [39].
This downstream application aims to evaluate the potential of Prithvi to estimate GPP at globally distributed eddy covariance sites. We utilize daily GPP data from eddy covariance flux towers, obtained from FLUXNET [40], AmeriFlux [^10] https://ameriflux.lbl.gov/data/flux-data-products/, and ICOS [41] networks. We use datasets processed by the ONEFLUX pipeline for high-quality and consistent data. The target variable for analysis is “GPP_NT_VUT_REF", which corresponds to GPP estimates derived from Net Ecosystem Exchange (NEE) measurements using the nighttime approach. We selected daily GPP data with at least 60% of high-quality hourly or half-hourly data for temporal aggregation. Metadata for flux tower sites, including latitude and longitude, is recorded to note site locations where satellite-based regressor variables are collected to prepare the dataset.
@@
These steps are replicated with all four variants of Prithvi models and we quantify the model’s skill in predicting GPP using the coefficient of determination, $R^{2}$ measure between the predicted and true GPP values. The $R^{2}$ performance in each train test split across all years is shown in Figure 16. Prithvi is also compared to a baseline model based on random forest, which uses the spatial average of HLS bands across each chip and the MERRA-2 data, representing the state-of-the-art approaches [39]. The baseline model also incorporated vegetation indices, including EVI, Normalized Difference Vegetation Index (NDVI), Normalized Difference Water Index (NDWI), Green Chlorophyll Index (GCI), near-infrared reflectance of vegetation (NIRv), and kernel NDVI. Results in Table 13 show that Prithvi-EO-2.0 models performed better than the random forest without vegetation indexes with up-to 20% improvement on $R^{2}$. Furthermore, the Prithvi-EO-2.0-600M version with temporal and location embeddings consistently outperformed the random forest including the vegetation indexes despite not having this information in its input. The proposed fine-tuning approach using Prithvi-EO-2.0 and MERRA-2 data shows promise for expanding the study to additional flux tower sites and extrapolating GPP beyond the flux tower network to derive a key climatic variable.